# Catalyst: Out-of-Distribution Detection via Elastic Scaling

### v1

@@ Abstract
Out-of-distribution (OOD) detection is critical for the safe deployment of deep neural networks. State-of-the-art post-hoc methods typically derive OOD scores from the output logits or penultimate feature vector obtained via global average pooling (GAP). We contend that this exclusive reliance on the logit or feature vector discards a rich, complementary signal: the raw channel-wise statistics of the pre-pooling feature map lost in GAP. In this paper, we introduce Catalyst, a post-hoc framework that exploits these under-explored signals. Catalyst computes an input-dependent scaling factor ($\gamma$) on-the-fly from these raw statistics (e.g., mean, standard deviation, and maximum activation). This $\gamma$ is then fused with the existing baseline score, multiplicatively modulating it – an “elastic scaling” – to push the ID and OOD distributions further apart. We demonstrate Catalyst is a generalizable framework: it seamlessly integrates with logit-based methods (e.g., Energy, ReAct, SCALE) and also provides a significant boost to distance-based detectors like KNN. As a result, Catalyst achieves substantial and consistent performance gains, reducing the average False Positive Rate by 32.87% on CIFAR-10 (ResNet-18), 27.94% on CIFAR-100 (ResNet-18), and 22.25% on ImageNet (ResNet-50). Our results highlight the untapped potential of pre-pooling statistics and demonstrate that Catalyst is complementary to existing OOD detection approaches. Our code is available here: https://github.com/epsilon-2007/Catalyst
@@ 1 Introduction
A deep neural network deployed in real-world environments will inevitably encounter out-of-distribution (OOD) samples – inputs drawn from novel contexts whose class labels are disjoint from the training distribution, referred as in-distribution (ID) data. Unlike ID samples that the model was trained on, these OOD instances should not be confidently classified but should be detected and flagged for human review. Robust OOD detection is particularly crucial for safety-critical applications where erroneous predictions can have severe consequences, e.g., in medical diagnosis [63, 52] or autonomous driving [9] systems.
Early method to OOD detection primarily focused on designing scoring functions to distinguish ID from OOD samples. The seminal work [16] proposed using the maximum softmax probability (MSP) as a confidence measure, based on observation that OOD samples yield lower softmax scores. However, subsequent studies [48, 15] revealed a critical flaw: neural networks often produce overconfident softmax predictions even for far-OOD inputs, rendering MSP unreliable. To address this, Energy [36] introduced the energy-based score, which maps inputs to a scalar value such that ID samples yield lower energy than OOD samples. This score provided a more robust uncertainty measure, inspiring a series of improvements aimed at enhancing ID-OOD separability. Recent advances have focused on post-hoc activation manipulation to amplify this separation. Notable methods includes ReAct [54], ASH [5], SCALE [67] achieving state-of-the-art performance.
@@
These methods share a common paradigm: they derive their scores using the penultimate feature vector (generally obtained via Global Average Pooling (GAP)) as their foundational input. These techniques process this feature vector to derive energy-based scores [36, 54, 67] or distance-based scores [56, 11]. We contend that exclusive reliance on feature vector creates an information bottleneck, as it discards complementary signal: the raw channel-wise statistics of the pre-pooling feature map – which could otherwise be used in tandem with existing methods for improved OOD detection.
@@
The existing methods have under-explored these distinctive statistical information. The exclusive reliance derived score from the output logit [16, 32, 36, 54, 5, 67] : discards potent raw cues (e.g., standard deviation, maximum) and fails to leverage independent discriminative power of raw mean statistics. To address this critical limitation, we propose Catalyst, a simple yet powerful framework that computes an input-dependent scaling factor ($\gamma$) designed to be fused in tandem with an existing scoring function. This scaling factor is computed on-the-fly, leveraging these distribution-sensitive cues embedded in the pre-pooled activation maps. Catalyst is designed to integrate seamlessly with established approaches while significantly improving their ability to distinguish between ID and OOD data. Our key contributions are:
@@
    Catalyst, a post-hoc OOD detection complementary framework that leverages pre-pooling channel-wise statistics to augment existing methods, generalizing across architectures like ResNet, DenseNet, and MobileNet.
@@
    An in-depth statistical analysis (Appendix B) and extensive ablation studies (Section 5) validate our design choices.
@@ 3.1 Computing the Scaling Factor $\gamma$
To compute scaling factor $\gamma$ we consider a trained DNN $\theta:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps an input $\mathbf{x}\in\mathbb{R}^{d}$ to a logit vector $f(\mathbf{x})\in\mathbb{R}^{C}$, where $C=|\mathcal{Y}|$ denotes the number of classes. The network’s penultimate layer produces a feature vector $h(\mathbf{x})\in\mathbb{R}^{n}$ by applying GAP operation to the activation map $g(\mathbf{x})\in\mathbb{R}^{n\times k\times k}$. Here, $n$ is the number of channels, and each channel has spatial resolution $k\times k$. A weight matrix $\mathbf{W}\in\mathbb{R}^{n\times C}$ projects $h(\mathbf{x})$ to the final logit vector.
In this work, we deliberately focus on this activation map $g(\mathbf{x})$ as our source of statistics. As we will empirically demonstrate in our ablation study (Section 5.1, Appendix H), this specific layer provides the most potent and reliable discriminative information cues for $\gamma$. The earlier layers provides less informative signal with high ID/OOD overlap.
@@
$\bar{f}(\mathbf{x})=\min(f(\mathbf{x}),c)\vskip-5.69054pt$ (3)
@@
where $f(\mathbf{x})\in\{h(\mathbf{x}),\sigma(\mathbf{x}),m(\mathbf{x})\}$ and the subscript $i$ denotes the $i$-th channel. The selection of this clipping threshold $c$ is discussed in Section 4.5 and detailed in Appendix G.
While we primarily focus on $\mu(\mathbf{x}),\sigma(\mathbf{x})$ and $m(\mathbf{x})$, our framework readily accommodates other channel-wise statistics derived from $g(\mathbf{x})$, such as entropy and median. We provide a detailed comparative analysis of these alternatives in our ablation study (Section 5.3 and Appendix J). This analysis provides justification for our design, validating our focus on $\mu(\mathbf{x}),\sigma(\mathbf{x})$, and $m(\mathbf{x})$ as the most robust and generalizable set of statistics for computing $\gamma$.
@@ 3.2 Elastic Scaling of the Score
To create a more discriminative score, we dynamically re-calibrate the baseline score $S(\mathbf{x};\theta)$ using scaling factor $\gamma$. We explore the two most direct fusion strategies: multiplicative and additive. We term the multiplicative strategy “Elastic Scaling” because it truly scales (i.e., multiplies) the baseline score, elastically stretching or shrinking it based on the $\gamma$ factor. The additive approach, in contrast, is a simple offset or shift, not a scaling. These are defined in Equation 5:
@@
While our analysis in Section 5.2 shows that both strategies can achieve similar peak performance, we adopt multiplicative fusion (Eq. 5a) as our primary framework. This choice is not arbitrary, as we demonstrate that the additive approach, while effective, is operationally fragile due to its hyperparameter sensitivity. The multiplicative fusion provides not only competitive performance but also the practical robustness and stability required of a general-purpose method. Therefore, in the remainder of this paper, we will refer multiplicative fusion as Elastic Scaling. This final re-calibrated score is subsequently used in the decision rule defined in Equation 2 to classify as ID or OOD.
@@ 3.3 Generalizability of Catalyst
While our analysis primarily focuses on energy-based scoring (given its role as the foundation for competitive methods like ReAct and SCALE), Catalyst is a general framework. It can be seamlessly integrated with other scoring functions such as MSP [16], ODIN [32], and KNN [56, 11] – by replacing the baseline score $S(\mathbf{x};\theta)$ in equation (Eq. 5a) with the alternate score.
@@ 4 Experiments
In this section, we evaluate the efficacy of Catalyst across diverse OOD datasets. We begin with an depth empirical analysis on the standard CIFAR benchmarks. Subsequently, we scale our evaluation to a large-scale OOD detection setting using ImageNet, demonstrating Catalyst’s versatility and robustness. Our evaluation is conducted without assuming the availability of an OOD validation set and incorporates wide range of OOD datasets to provide a realistic assessment of Catalyst.
We use the foundational Energy score [36] as our default baseline. For brevity, when Catalyst is applied to Energy, we simply denote it as Catalyst. When applying Catalyst to other baselines, we state it explicitly (e.g., Catalyst + ReAct)
We ensure a fair and direct comparison against prior work. As the architectures used in our evaluation (e.g., ResNet-18 for the CIFAR benchmarks; ResNet-34 and DenseNet-121 for ImageNet) were not included in the foundational baselines like ReAct, DICE, ASH, and SCALE, we undertook a rigorous re-evaluation of these methods. We carefully followed the official hyperparameter selection protocols and open-sourced implementations from their respective papers to ensure the integrity of our comparisons.
@@ 4.1 CIFAR Evaluation
Results. Table 1 summarizes our results on the CIFAR benchmarks. The table clearly shows the two key benefits(1) Catalyst (e.g, ${\texttt{Catalyst}}(m)$) significantly outperforms the standard energy score baseline, proving the inherent value of scaling factor. (2) When composed with ReAct, Catalyst establishes a new benchmark. For instance, on CIFAR-10, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces FPR95 by 32.87%, 28.10% with ResNet-18 and DenseNet-101 respectively. On CIFAR-100, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces FPR95 by 27.94% and 18.99% with ResNet-18 and DenseNet-101 respectively. The detailed per-dataset results are provided in Appendix E.3.
Near-OOD Evaluation. We also evaluate Catalyst on the challenging near-OOD task of distinguishing CIFAR-10 from CIFAR-100 [11]. Catalyst provides a consistent performance boost when applied in tandem with baselines, demonstrating its robustness. For brevity, detailed results are available in Appendix E.1.
@@ 4.2 ImageNet Evaluation
Experimental Setup. To assess scalability and performance in a more realistic setting, we evaluate on the ImageNet-1k benchmark. We use four OOD datasets: iNaturalist [60], SUN [66], Places365 [73], and Textures [3]. These datasets are carefully curated to avoid class overlap with ImageNet, while spanning distinct semantic domains to rigorously assess generalization performance [36, 54].
Our evaluation showcases broad architectural robustness by using pre-trained ResNet-34, ResNet-50, DenseNet-121, and MobileNet-v2. Since foundational baselines (e.g., ReAct, SCALE) did not originally report results on all of these architectures (such as ResNet-34 and DenseNet-121), we undertook a rigorous re-evaluation of all methods.
Results. Table 2 shows that Catalyst yields consistent improvements at ImageNet scale. Compared to energy score, ${\texttt{Catalyst}}(m)$ reduces FPR95 by 44.35%, 47.99%, 43.69%, and 21.23% using ResNet-34, ResNet-50, MobileNet-v2, and DenseNet-121 architectures respectively. The most significant gains are achieved when composing Catalyst with existing foundational methods like ReAct. Specifically, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ improves FPR95 by 25.39%, 19.41%, 5.57% and 12.62% compared to previous best results using ResNet-34, ResNet-50, MobileNet-v2, and DenseNet-121 respectively. These results validate that the principles of Catalyst are effective in complex, large-scale scenarios and across diverse architectural families. The performance boost confirms that our pre-pooling scaling factor provides significant discriminative information that is complementary to both foundational and existing competitive techniques. The detailed per-dataset results are in Appendix E.2.
@@ 4.4 Generalizability to Distance-Based Methods
To validate Catalyst as a general-purpose framework, we test its synergy with a distance-based K-Nearest Neighbors (KNN) [56, 11] OOD detector, applied to our standard pre-trained models.
The results in Tables 3 and 4 confirm our hypothesis, showing that Catalyst provides significant improvement over the KNN baseline across all benchmarks. For instance, on CIFAR-100 (ResNet-18), ${\texttt{Catalyst}}(m)$ achieves a 43.84% reduction. Similarly on, large-scale ImageNet benchmark, where ${\texttt{Catalyst}}(\mu)$ on a ResNet-50 results in a 52.13% reduction in average FPR95. These performance boost highlight that Catalyst is a general-purpose modulator, providing complementary information for both logit and distance based methods, making it a true plug-and-play framework. The full experimental setup and detailed per-dataset results are in Appendix D.
@@ 4.5 Hyperparameter Selection
The clipping threshold $\boldsymbol{c}$ (Eq. 3) is crucial for enhanced performance, as it must be set to optimally distinguish ID from OOD data. Analogous to ReAct [54], we set $\boldsymbol{c}$ to the $p$-th percentile of the ID activation distribution. The choice of this percentile $p$ is the key hyperparameter to be tuned. To demonstrate its sensitivity, we summarize the OOD detection performance of ${\texttt{Catalyst}}(m)$ in Table 3, varying $p$ from 10 to 100 at 5-point intervals. To this end, we follow established protocols [55, 54] and create a proxy OOD validation set, generated by adding pixel-wise Gaussian noise to images from the ID validation set. We then select the percentile $p$ that yields the best OOD separation on this proxy task. This two-step procedure – using a percentile for the mechanism and a proxy set for tuning – is a robust tuning strategy grounded in prior work. The specific details and the selected $p$ values are provided in Appendix G.

### v2

@@ Abstract
Out-of-distribution (OOD) detection is critical for the safe deployment of deep neural networks. State-of-the-art post-hoc methods typically derive OOD scores from the output logits or penultimate feature vector obtained via global average pooling (GAP). We contend that this exclusive reliance on the logit or feature vector discards a rich, complementary signal: the raw channel-wise statistics of the pre-pooling feature map lost in GAP. In this paper, we introduce Catalyst, a post-hoc framework that exploits these under-explored signals. Catalyst computes an input-dependent scaling factor ($\gamma$) on-the-fly from these raw statistics (e.g., mean, standard deviation, and maximum activation). This $\gamma$ is then fused with the existing baseline score, multiplicatively modulating it – an “elastic scaling” – to push the ID and OOD distributions further apart. We demonstrate Catalyst is a generalizable framework: it seamlessly integrates with logit-based methods (e.g., Energy, ReAct, SCALE) and also provides a significant boost to distance-based detectors like KNN. As a result, Catalyst achieves substantial and consistent performance gains, reducing the average False Positive Rate by 32.87% on CIFAR-10 (ResNet-18), 27.94% on CIFAR-100 (ResNet-18), and 22.25% on ImageNet (ResNet-50). Our results highlight the untapped potential of pre-pooling statistics and demonstrate that Catalyst is complementary to existing OOD detection approaches. Our code is available here: https://github.com/bingabid/Catalyst
@@ 1 Introduction
A deep neural network deployed in real-world environments will inevitably encounter out-of-distribution (OOD) samples drawn from novel contexts whose class labels are disjoint from the training distribution, referred as in-distribution (ID) data. Unlike ID samples that the model was trained on, these OOD instances should not be confidently classified but be detected and flagged for human review. Robust OOD detection is particularly crucial for safety-critical applications where erroneous predictions can have severe consequences, e.g., in medical diagnosis [63, 52] or autonomous driving [9] systems.
Early methods to OOD detection primarily focused on designing scoring functions to distinguish ID from OOD samples. The seminal work [16] proposed using the maximum softmax probability (MSP) as a confidence measure, based on observation that OOD samples yield lower softmax scores. However, subsequent studies [48, 15] revealed a critical flaw: neural networks often produce overconfident softmax predictions even for far-OOD inputs, rendering MSP unreliable. To address this, Energy [36] introduced the energy-based score, which maps inputs to a scalar value such that ID samples yield lower energy than OOD samples. This score provided a more robust uncertainty measure, inspiring a series of improvements aimed at enhancing ID-OOD separability. Recent advances have focused on post-hoc activation manipulation to amplify this separation. Notable methods include ReAct [54], DICE [55], ASH [5], SCALE [67] achieving state-of-the-art performance.
@@
These methods share a common paradigm: they derive their scores using the penultimate feature vector (generally obtained via GAP) as their foundational input. These techniques process this feature vector to derive energy-based scores [36, 54, 67] or distance-based scores [56, 11]. We contend that exclusive reliance on the feature vector creates an information bottleneck, as it discards complementary signals, namely the raw channel-wise statistics of the pre-pooling feature map, which could otherwise be used in tandem with existing methods for improved OOD detection.
@@
The existing methods have under-explored these distinctive statistical information. The approach exclusively relies on a score derived from the output logits [16, 32, 36, 54, 5, 67]: discards potent raw cues (e.g., standard deviation, maximum) and fails to leverage independent discriminative power of raw mean statistics. To address this critical limitation, we propose Catalyst, a simple yet powerful framework that computes an input-dependent scaling factor ($\gamma$) designed to be fused in tandem with an existing scoring function. This scaling factor is computed on-the-fly, leveraging these distribution-sensitive cues embedded in the pre-pooled activation maps. Catalyst is designed to integrate seamlessly with established approaches while significantly improving their ability to distinguish between ID and OOD data. Our key contributions are:
@@
    Catalyst, a complementary post-hoc OOD detection framework that leverages pre-pooling channel-wise statistics to augment existing methods, generalizing across architectures like ResNet, DenseNet, and MobileNet.
@@
    Statistical analysis (Appendix B) and extensive ablation studies (Section 5) validate our design choices.
@@ 3.1 Computing the Scaling Factor $\gamma$
To compute scaling factor $\gamma$, we consider a trained DNN $\theta:\mathbb{R}^{d}\rightarrow\mathbb{R}^{C}$ that maps an input $\mathbf{x}\in\mathbb{R}^{d}$ to a logit vector $f(\mathbf{x})\in\mathbb{R}^{C}$, where $C=|\mathcal{Y}|$ denotes the number of classes. The network’s penultimate layer produces a feature vector $h(\mathbf{x})\in\mathbb{R}^{n}$ by applying GAP operation to the activation map $g(\mathbf{x})\in\mathbb{R}^{n\times k\times k}$. Here, $n$ is the number of channels, and each channel has spatial resolution $k\times k$. A weight matrix $\mathbf{W}\in\mathbb{R}^{n\times C}$ projects $h(\mathbf{x})$ to the final logit vector.
In this work, we deliberately focus on this activation map $g(\mathbf{x})$ as our source of statistics. As we will empirically demonstrate in our ablation study (Section 5.1, Appendix H), this specific layer provides the most potent and reliable discriminative information cues for $\gamma$. The earlier layers provide less informative signal with high ID/OOD overlap.
@@
$\bar{f}(\mathbf{x})=\min(f(\mathbf{x}),c)\vskip-2.84526pt$ (3)
@@
where the subscript $i$ denotes the $i$-th channel. The selection of this clipping threshold $c$ is discussed in Section 4.5 and detailed in Appendix G.
While we primarily focus on $\mu(\mathbf{x}),\sigma(\mathbf{x})$ and $m(\mathbf{x})$, our framework readily accommodates other channel-wise statistics derived from $g(\mathbf{x})$, such as entropy and median. We provide a detailed comparative analysis of these cues in our ablation study (Section 5.3; Appendix J), which justifies our design and validates our focus on $\mu(\mathbf{x}),\sigma(\mathbf{x})$, and $m(\mathbf{x})$ as robust and generalizable set of statistics for computing $\gamma$.
@@ 3.2 Elastic Scaling of the Score
To create a more discriminative score, we dynamically re-calibrate the baseline score $S(\mathbf{x};\theta)$ using scaling factor $\gamma$. We explore the two fusion strategies: multiplicative and additive. We term the multiplicative strategy “Elastic Scaling” because it truly scales (i.e., multiplies) the baseline score, elastically stretching or shrinking it based on the $\gamma$. The additive approach, in contrast, is a simple offset or shift, not a scaling. These are defined in Equation 5:
@@
While our analysis in Section 5.2 shows that both strategies can achieve similar peak performance, we adopt multiplicative fusion (Eq. 5a) as our primary framework. This choice is not arbitrary, as we demonstrate that the additive approach, while effective, is operationally fragile due to its hyperparameter sensitivity. The multiplicative fusion provides not only competitive performance but also the practical robustness and stability required of a general-purpose usage. Therefore, in the remainder of this paper, we will refer multiplicative fusion as Elastic Scaling. This final re-calibrated score is subsequently used in the decision rule defined in Equation 2 to classify as ID or OOD.
@@ 3.3 Generalizability of Catalyst
While our analysis primarily focuses on energy-based scoring (given its primary role for competitive methods like ReAct and SCALE), Catalyst is a general framework. It can be seamlessly integrated with other scoring functions such as MSP [16], ODIN [32], and KNN [56, 11] – by replacing the baseline score $S(\mathbf{x};\theta)$ in equation (Eq. 5a) with the alternate score.
@@ 4 Experiments
In this section, we evaluate the efficacy of Catalyst across a diverse set of OOD datasets. We begin with an in-depth empirical analysis on standard CIFAR benchmarks. We then extend our evaluation to a large-scale OOD detection setting using ImageNet, demonstrating the versatility and robustness of Catalyst. Our evaluation does not assume the availability of an OOD validation set and incorporates a wide range of OOD datasets to provide a realistic assessment of Catalyst.
We use the Energy score [36] as our default baseline. For brevity, when Catalyst is applied to Energy, we simply denote it as Catalyst. When applying Catalyst to other baselines or scoring functions, we state it explicitly (e.g., Catalyst + ReAct, Catalyst + KNN).
We ensure a fair and direct comparison against prior work. As the architectures used in our evaluation (e.g., ResNet-18 for the CIFAR benchmarks; ResNet-34 and DenseNet-121 for ImageNet) were not included in the primary baselines like ReAct, DICE, ASH, and SCALE, we undertook a rigorous re-evaluation of these methods. We carefully followed the official hyperparameter selection protocols and open-sourced implementations from their respective papers to ensure the integrity of our comparisons.
@@ 4.1 CIFAR Evaluation
Results. Table 1 summarizes our results on the CIFAR benchmarks. The table clearly shows the two key benefits(1) Catalyst (e.g., ${\texttt{Catalyst}}(m)$) significantly outperforms the standard energy score baseline, proving the inherent value of scaling factor. (2) When composed with ReAct, Catalyst establishes a new benchmark. For instance, on CIFAR-10, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces FPR95 by 32.87%, 28.10% with ResNet-18 and DenseNet-101 respectively. On CIFAR-100, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ reduces FPR95 by 27.94% and 18.99% with ResNet-18 and DenseNet-101 respectively. The detailed per-dataset results are provided in Appendix E.3.
Near-OOD Evaluation. We evaluate Catalyst on the challenging near-OOD task of distinguishing CIFAR-10 from CIFAR-100 [11]. While it yields marginal gains over SCALE, the improvements are less pronounced than in far-OOD settings, likely due to the high similarity of learned penultimate representations. Designing a more effective $\gamma$ for near-OOD settings remains an important direction for future work. Detailed results are provided in Appendix E.1.
@@ 4.2 ImageNet Evaluation
Experimental Setup. To assess scalability in a more realistic setting, we evaluate on the ImageNet-1k benchmark. We use four OOD datasets: iNaturalist [60], SUN [66], Places365 [73], and Textures [3]. These datasets are carefully curated to avoid class overlap with ImageNet, while spanning distinct semantic domains to rigorously assess generalization performance [36, 54].
Our evaluation showcases broad architectural robustness by using pre-trained ResNet-34, ResNet-50, DenseNet-121, and MobileNet-v2. Since primary baselines (e.g., ReAct, SCALE) did not originally report results on all of these architectures (such as ResNet-34 and DenseNet-121), we undertook a rigorous re-evaluation of all methods.
Results. Table 2 shows that Catalyst yields consistent improvements at ImageNet scale. Compared to energy score, ${\texttt{Catalyst}}(m)$ reduces FPR95 by 44.35%, 47.99%, 43.69%, and 21.23% using ResNet-34, ResNet-50, MobileNet-v2, and DenseNet-121 architectures respectively. The most significant gains are achieved when composing Catalyst with existing primary methods like ReAct. Specifically, ${\texttt{Catalyst}}(m)+\texttt{ReAct}$ improves FPR95 by 25.39%, 19.41%, 5.57% and 12.62% compared to previous best results using ResNet-34, ResNet-50, MobileNet-v2, and DenseNet-121 respectively. These results validate that the principles of Catalyst are effective in complex, large-scale datasets and across diverse architectural families. The performance boost confirms that the scaling factor $\gamma$ provides significant discriminative information that is complementary to existing competitive techniques. The detailed per-dataset results are in Appendix E.2.
@@ 4.4 Generalizability to Distance-Based Methods
To validate Catalyst as a general-purpose framework, we test its synergy with a distance-based K-Nearest Neighbors (KNN) [56, 11] OOD detector. The results in Tables 3 and 4 confirm our hypothesis, showing that Catalyst provides significant improvement over the KNN baseline across all benchmarks. For instance, on CIFAR-100 (ResNet-18), ${\texttt{Catalyst}}(m)$ achieves a 43.84% reduction. Similarly on, large-scale ImageNet benchmark, where ${\texttt{Catalyst}}(\mu)$ on a ResNet-50 results in a 52.13% reduction in average FPR95. These performance boost highlight that Catalyst is a general-purpose modulator, providing complementary information for both logit and distance based methods, making it a true plug-and-play framework. The full experimental setup and detailed per-dataset results are in Appendix D.
@@ 4.5 Hyperparameter Selection
The clipping threshold $\boldsymbol{c}$ (Eq. 3) is crucial for enhanced performance, as it must be set to optimally distinguish ID from OOD data. Analogous to ReAct [54], we set $\boldsymbol{c}$ to the $p$-th percentile of the ID activation distribution. The choice of this percentile $p$ is the key hyperparameter to be tuned. To demonstrate its sensitivity, we summarize the OOD detection performance of ${\texttt{Catalyst}}(m)$ in Figure 3, varying $p$ from 10 to 100 at 5-point intervals. To this end, we follow established protocols [55, 54] and create a proxy OOD validation set, generated by adding pixel-wise Gaussian noise to images from the ID validation set. We then select the percentile $p$ that yields the best OOD separation on this proxy task. This two-step procedure, which uses a percentile for the mechanism and a proxy set for tuning, is a robust tuning strategy grounded in prior work. The specific details and the selected $p$ values are provided in Appendix G.